5. Use the point-slope from to write an equation of the line the given properties.
slope -4;passes through (7,0)
Write the equation in slope-intercept form.
I really need help with slope and-intercept

Found 2 solutions by stanbon, Shana-D77:
Answer by stanbon(75887)

Use the point-slope from to write an equation of the line the given properties.
slope -4;passes through (7,0)
Write the equation in slope-intercept form.
I really need help with slope and-intercept
-------
Point-slope form: y-y1 = m(x-x1)
y - 0 = -4(x-7)
y = -4x+28
----------
Slope-intercept form.
y = -4x+28
slope = -4
y-intercept = f(0) = +28
==============================
Cheers,
Stan H.
==============

Answer by Shana-D77(132)

When you're given a slope and a point, it's best to use the point-slope form of the equation. "The" equation means the linear (straight line) equation. The point slope form is:
y = y1 = m(x - x1).
((side note: this is actually the slope formula, just rearranged.))
y = mx + b only works if you have b (the y-intercept). A y intercept always has an x value of 0. Examples: (0, 4), (0, -2), etc). Your teacher gave you the x-intercept. I know this because (7, 0) has a y value of 0, not an x value of 0.
So,
slope (m) = -4
point (7, 0)
y - y1 = m(x - x1)
(The y and the x always stay as y and x in this form)
Plug in what your teacher gave you. y1 = 0 and x1 = 7 (from (7,0)), and slope (m) = -4.
y - 0 = -4(x - 7)
y - 0 = -4x + 28 (distributed the -4)
y = -4x + 28